what is 8s+9=7s-21
Answer by hyh_123(8)

8s+9=7s-21
8s-7s+9=-21 (same terms on one side)
s =-21-9 (negative sign get added together)
s =-30
CHECK
8(-30)+9=7(-30)-21
-210+9=-210-21
-231=-231 *